OpenStudy (anonymous):

integral cosθ/sin^2 θ+4sinθ−5 dθ

OpenStudy (anonymous):

\[\int\left(\frac{\cos{\theta}}{\sin^2\theta} + 4\sin\theta - 5\right)d\theta\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no it's \[\int\limits \cos \theta / \sin^2 \theta + 4 \sin \theta - 5\]